# Match-and-Fuse: Consistent Generation from Unstructured Image Sets

### v1

@@ Abstract
We present Match-and-Fuse – a zero-shot, training-free method for consistent controlled generation of unstructured image sets – collections that share a common visual element, yet differ in viewpoint, time of capture, and surrounding content. Unlike existing methods that operate on individual images or densely sampled videos, our framework performs set-to-set generation: given a source set and user prompts, it produces a new set that preserves cross-image consistency of shared content. Our key idea is to model the task as a graph, where each node corresponds to an image and each edge triggers a joint generation of image pairs. This formulation consolidates all pairwise generations into a unified framework, enforcing their local consistency while ensuring global coherence across the entire set. This is achieved by fusing internal features across image pairs, guided by dense input correspondences, without requiring masks or manual supervision. It also allows us to leverage an emergent prior in text‑to‑image models that encourages coherent generation when multiple views share a single canvas. Match-and-Fuse achieves state-of-the-art consistency and visual quality, and unlocks new capabilities for content creation from image collections. Code and data will be released at our project page.
@@ 2 Related work
Text-to-image controlled generation.  T2I models [rombach2022highresolutionimagesynthesislatent, saharia2022photorealistictexttoimagediffusionmodels, shi2020improvingimagecaptioningbetter, fluxcontrolnet2024, esser2024scalingrectifiedflowtransformers] have made a remarkable progress. Numerous methods extended T2I models to go beyond input text and condition the generation on various signals, such as style (e.g., [ye2023ip-adapter, mou2023t2i]) or various spatial controls: pose, depth, or edge maps [zhang2023addingconditionalcontroltexttoimage]. A growing line of work tackles image editing rather than conditioning, including trainable editing models [brooks2022instructpix2pix, labs2025flux1kontextflowmatching, google2025nanobanana] that learn to modify an image according to text or paired examples, as well as training-free editing approaches [Tumanyan_2023_CVPR, hertz2022prompt, zhu2025kv] that manipulate diffusion trajectories at inference time. However, all these methods operate on individual images and lack the set-level consistency mechanisms required for our task.
@@
Storyboard generation.  This line of work aims to produce image sequences depicting consistent recurring characters across frames [tewel2024training, avrahami2024chosen, ye2023ip-adapter, garibi2025tokenverseversatilemulticonceptpersonalization], or emphasizing narrative coherence across prompts [zhou2024storydiffusion, Liu_2024_CVPR, he2024dreamstory, lhhuang2024iclora]. Zero-shot methods manipulate attention features through injected correspondences or masks [tewel2024training, zhou2024storydiffusion], whereas training-based approaches fine-tune diffusion models for improved consistency [he2024dreamstory, Liu_2024_CVPR, lhhuang2024iclora]. Aside from the limited pose control offered by [tewel2024training], these models rely solely on text and lack spatial control mechanisms necessary for set-to-set generation.
Consistent image-set generation.  Related to our work, Edicho [bai2024edicho] addresses pairwise consistent editing, first modifying one image and then transferring the edit to its partner using 2D correspondences to warp intermediate features. Their method is similar to ours in leveraging explicit matches and operating on unstructured image sets, but it remains strictly pairwise: edits are propagated only from a single reference image, causing coherence to degrade for views farther from that anchor. In contrast, [lhhuang2024iclora] enforces consistency by generating all images simultaneously on a single canvas, where a LoRA module [hu2022lora] is fine-tuned on the multi-image prompt. This design ties the method to a fixed grid, limiting its scalability to larger image sets.
3D and video-based approaches. Distinct from these methods, 3D editing techniques [instructnerf2023, vicanerf2023, wang2024view, gaussctrl2024] enable manipulation of static multi-view sets through rendered views but operate on 3D representations rather than directly editing 2D images. This imposes restrictive assumptions that limit their applicability to image sets with unknown camera parameters, articulated objects, and varying backgrounds. Similarly, video editing and generation methods (e.g., [tokenflow2023, yuan2024instructvideo, wang2024cove]) are unsuitable for our task, as they assume temporal continuity and rely on motion coherence that does not hold across unordered image sets.
@@ 3 Method
We build on a pre-trained, frozen, depth-conditioned T2I model. Although designed for single-image generation, these models have been shown to produce image grids when prompted with joint layouts (e.g., “Side-by-side views of…”), establishing cross-image relationships as demonstrated in recent work [lhhuang2024iclora, shin2024diptychprompting]. However, this emerged capability, which we refer to as the grid prior, exhibits several key limitations: (i) it provides only partial consistency in appearance, shape, and semantics (Fig. 2b); (ii) the consistency deteriorates rapidly as more images are composed (Fig. 2b); and (iii) generating a single canvas is bounded by the model’s native resolution, limiting scalability.
Our method leverages the grid prior while overcoming its core limitations (Fig. 2d). Specifically, we model the image set as a Pairwise Consistency Graph, comprising all possible two-image grid generations (Sec. 3.2). This allows us to exploit the strong inductive bias of the grid prior, while eliminating its scale limitation. To enhance visual consistency both within each image grid and across grids, we perform joint feature manipulation across all pairwise generations. To this end, we utilize dense 2D correspondences from the source set to automatically identify shared regions – without requiring object masks – and enforce fine-grained alignment. We find that feature-space similarity along these matches correlates strongly with visual coherence, motivating the use of Multiview Feature Fusion. We further refine details via Feature Guidance, using a feature-matching objective (Sec. 3.4). Our full pipeline is illustrated in Fig. 3.
@@ 3.2 Pairwise Consistency Graph
Since every node $z^{t}_{i}$ participates in multiple adjacent edges, a graph denoising step yields multiple versions of $z^{t-1}_{i|ij}$. We thus consolidate pairwise latents $\{z^{t-1}_{ij}\}_{e\in E}$ back into image latents $\{z^{t-1}_{i}\}_{i=1}^{N}$ by extracting and averaging all $z^{t-1}_{i|ij}$, following [bar2023multidiffusion]. In practice, full graph connectivity is not required, as analyzed in Sec. 4.4.
@@ 3.3 Multiview Feature Fusion
Matches between each image pair are defined as a partial coordinate mapping in feature space, $M_{ij}:\mathcal{C}_{i}\!\to\!\mathcal{C}_{j}$, where a coordinate $\mathbf{c}\in I_{i}$ corresponds to $\mathbf{c}^{\prime}=M_{ij}(\mathbf{c})\in I_{j}$ and $\mathcal{C}_{i}$ contains all matched points (Fig. 4b). We first address pairwise consistency by considering a two-node graph with a single edge. Let $\mathbf{f}_{ij}\in\mathbb{R}^{H\times 2W\times D}$ be a feature map from a model’s forward pass on a grid, interpreted as $\mathbf{f}_{ij}=\operatorname{concat}(\mathbf{f}_{i},\mathbf{f}_{j})$, with $\mathbf{f}_{i},\mathbf{f}_{j}\in\mathbb{R}^{H\times W\times D}$. Given a pooling operator $\mathbf{f}[\mathbf{c}]$ extracting the vector at $\mathbf{c}$, Multiview Feature Fusion (MFF) fuses matched coordinates:
@@ 3.4 Feature Guidance
Figure 5: Matched feature similarity vs. visual consistency. We consider increasingly consistent generation (left to right): random images $\to$ adding descriptive prompts $\to$ adding control signals $\to$ generating in a grid $\to$ DDIM [song2020denoising] inversion which reconstructs fully consistent source images. Keys and values differ in scale but follow the same pattern: cosine similarity at matched locations rises with consistency. Dashed lines show the baseline all-to-all similarity of feature maps. Points are averaged over 40 image pairs, all correspondences, blocks, and timesteps.
@@
To align small details, we further perform feature guidance [epstein2023selfguidance, chefer2023attendandexciteattentionbasedsemanticguidance, shi2023dragdiffusion], completing each generation step (Fig. 3, right). Specifically, we define an optimization objective over all edges as the distance between matched features, and refine all node latents via gradient descent with $\nabla_{z^{t-1}_{i}}L_{\mathrm{guide}}$ of:
@@ 4 Experiments
Figure 9: Qualitative comparisons. Match-and-Fuse (Ours) generates image sets with the highest consistency compared to our baselines FLUX [flux2024], FLUX Kontext [labs2025flux1kontextflowmatching], IC-LoRA [lhhuang2024iclora], and Edicho [bai2024edicho]. See Sec. 4.3 for comments on each result and SM for full image sets.
Figure 10: Extended Applications. Our method generalizes across various settings, enabling (a) consistent story generation from sketches and (b) localized editing omitting $\mathcal{P}^{theme}$ via FlowEdit [kulikov2024flowedit] integration. See full sets of results in SM.
We extensively evaluate Match-and-Fuse. Our implementation builds on FLUX [flux2024] with depth-conditioning [fluxcontrolnet2024], and uses RoMA [edstedt2023romarobustdensefeature] for matching. Full details are in SM.
@@ 4.1 Evaluation Setup
Evaluating our task requires image sets that share content across diverse scenarios, along with a metric that captures fine-grained cross-image consistency. Since no existing benchmark or metric supports this setting, we introduce a setup tailored to consistent image-set generation.
Evaluation set.  We curate a benchmark with 400 edits for 149 diverse distinct image sets, 3–15 images in each, combining all sets from [kumari2022customdiffusion] and [ruiz2023dreambooth] with frames sampled from 3D datasets [instructnerf2023, barron2021mipnerf, li2022neural3dvideosynthesis], keyframes of a few publicly sourced videos, and sketched storyboards generated by ChatGPT. Data will be publicly released.
@@
Metrics.  We quantitatively evaluate our method, beginning with fine-grained cross-image consistency. Previous work relies on global visual-similarity metrics to assess set consistency—for example, [tewel2024training] uses DreamSim [fu2023dreamsim] on object-masked regions. However, as our user study (Tab. 2b) shows, such global metrics are not ideal. To address this, we introduce DINO-MatchSim. For each image pair, we compute patch-level nearest-neighbor correspondences $\textit{NN}_{ij}$ between DINOv3 [simeoni2025dinov3] feature maps from the source images, and measure similarity at the corresponding output locations:
@@
where $\tilde{F}_{i}$ denotes the $i^{\text{th}}$ output feature map. The average similarity across all image pairs is DINO-MatchSim. For prompt adherence, we report the average CLIP score over the set. To assess human preference, we conduct a large-scale user study on a random subset of the benchmark, comparing Match-and-Fuse to a baseline in a randomized two-alternative forced-choice (2AFC) format. Finally, we include a VLM-based evaluation using GPT-5 [openai2025gpt5], which received the same content as the human raters.
@@ 4.3 Comparisons
Since no existing method is designed for our set-to-set generation task, we compare Match-and-Fuse to the closest baselines: (1) IC-LoRA [lhhuang2024iclora], which fine-tunes LoRA modules for specific context types to produce text-driven image sets with shared elements. We use their most relevant public checkpoint and add spatial ControlNet conditioning for comparability. (2) FLUX Kontext [labs2025flux1kontextflowmatching], an image-to-image editing model. While it provides consistent edits over multiple turns on a single image, it lacks an explicit mechanism for achieving consistency across a single-turn image set editing. (3) FLUX [flux2024] with ControlNet [fluxcontrolnet2024] conditioning, which serves as a baseline for assessing independent generation using the same base model as our method. (4) Edicho [bai2024edicho], the closest baseline to our method, combines depth-conditioned generation with explicit correspondences, but is limited to an image pair editing through a single reference. Additional implementation details are in Sec. A.2.
Figure 9 shows representative qualitative results. FLUX [flux2024] produces inconsistencies in both coarse and fine details. FLUX Kontext [labs2025flux1kontextflowmatching] has low prompt adherence (e.g., a toy remains animal-like), and often distorts object structure (e.g. vase, dog). IC-LoRA [lhhuang2024iclora] achieves partial consistency, with coherence often restricted to subsets (vase). Its realism and fidelity are notably lower than FLUX. Edicho [bai2024edicho] performs best among baselines but still shows noticeable inconsistencies (dog), as its one-to-all warping enforces consistency only with the first image, making the choice of an anchor ambiguous and degrading coherence across views. In contrast, Match-and-Fuse (bottom row) maintains high image quality while substantially improving structural and fine-grained consistency. We additionally compare to the closed-source Nano Banana API [google2025nanobanana] and find that it does not solve our task (SM).
@@ 4.5 Extended Applications
As shown in Fig. 10a, our method generalizes to sketched source sets, enabling consistent visualization of storyboards. In Fig. 10b, we further demonstrate localized editing by combining our method with FlowEdit [kulikov2024flowedit]. We provide default integration parameters in the SM; however, as is a known limitation of FlowEdit, achieving the desired balance between structure preservation and appearance change often requires mild per-edit hyperparameter tuning. Automating this selection is left for future work.
@@ A.1 Method
Preprocessing. Prompt construction with a VLM (Sec. 3.1) is used as an automated, scalable alternative to manual prompting. We instruct GPT-4o [openai2024gpt4o] via the system prompt (Fig. A.3) to generate $P^{non-shared}_{i}$ and enrich $\mathcal{P}^{\text{shared}}$ with creative artistic details (e.g., colors, textures). Figure A.1 exemplifies VLM input and output. Source 2D correspondences and shared content are identified using RoMA [edstedt2023romarobustdensefeature] predictions filtered by a confidence threshold $\operatorname{c}>0.05$.
General. All results are generated using FLUX [flux2024] with $t=25\dots 0$ steps. We apply ControlNet [fluxcontrolnet2024] complementing DepthAnythingV2 [yang2024depthv2] depth with HED edge maps [xie2015holisticallynestededgedetection] with equal weights, and linearly anneal control strength $w_{\text{ctrl}}^{t}=\text{linear\_anneal}(1,0,t)$. This enforces correct layout while preventing over-constraining control.
@@
Localized editing. FlowEdit [kulikov2024flowedit] is an inversion-free method that enables background-preserving editing by traversing a path between source and target image distributions with $v$-predictions $V=V^{tar}(z^{tar})-V^{src}(z^{src})$. To incorporate it into our method, we redefine $V_{ij}$ implied in $\operatorname{denoise}_{\operatorname{MFF}}(\cdot)$ in Eq. 1 as $V_{ij}=V^{tar}_{ij}-V^{src}_{ij}$, where
@@
Similar to [meng2022sdeditguidedimagesynthesis], FlowEdit relies on a single starting timestep parameter $n_{\text{max}}$ to balance structure preservation and appearance deviation. We ease this trade-off by adopting spatial control with an earlier start step. The used hyperparameters are: $T=25,n_{\text{max}}=24,n_{\text{min}}=10,w_{\text{ctrl}}=0.5,t^{MFF}_{\text{d}}>20$. Feature Guidance is omitted.

### v2

@@ Abstract
We present Match-and-Fuse – a zero-shot, training-free method for consistent controlled generation of unstructured image sets – collections that share a common visual element, yet differ in viewpoint, time of capture, and surrounding content. Unlike existing methods that operate on individual images or densely sampled videos, our framework performs set-to-set generation: given a source set and user prompts, it produces a new set that preserves cross-image consistency of shared content. Our key idea is to model the task as a graph, where each node corresponds to an image and each edge triggers a joint generation of image pairs. This formulation consolidates all pairwise generations into a unified framework, enforcing local consistency while ensuring global coherence across the entire set. This is achieved by fusing internal features across image pairs, guided by dense input correspondences, without requiring masks or manual supervision, and by leveraging an emergent prior in text-to-image models that encourages coherent generation when multiple views share a single canvas. Match-and-Fuse achieves state-of-the-art consistency and visual quality, and unlocks new capabilities for content creation from image collections. Code and data: project page.
@@ 2 Related work
Text-to-image controlled generation.  T2I models [33, 35, 37, 49, 14] have made a remarkable progress. Numerous methods extended T2I models to go beyond input text and condition the generation on various signals, such as style (e.g., [51, 29]) or various spatial controls: pose, depth, or edge maps [54]. A growing line of work tackles image editing rather than conditioning, including trainable editing models [7, 25, 10] that learn to modify an image according to text or paired examples, as well as training-free editing approaches [43, 20, 56] that manipulate diffusion trajectories at inference time. However, all these methods operate on individual images and lack the set-level consistency mechanisms required for our task.
@@
Storyboard generation.  This line of work aims to produce image sequences depicting consistent recurring characters across frames [42, 2, 51, 16], or emphasizing narrative coherence across prompts [55, 27, 19, 22]. Zero-shot methods manipulate attention features through injected correspondences or masks [42, 55], whereas training-based approaches fine-tune diffusion models for improved consistency [19, 27, 22]. Aside from the limited pose control offered by [42], these models rely solely on text and lack spatial control mechanisms necessary for set-to-set generation.
Consistent image-set generation.  Related to our work, Edicho [3] addresses pairwise consistent editing, first modifying one image and then transferring the edit to its partner using 2D correspondences to warp intermediate features. Their method is similar to ours in leveraging explicit matches and operating on unstructured image sets, but it remains strictly pairwise: edits are propagated only from a single reference image, causing coherence to degrade for views farther from that anchor. In contrast, [22] enforces consistency by generating all images simultaneously on a single canvas, where a LoRA module [21] is fine-tuned on the multi-image prompt. This design ties the method to a fixed grid, limiting its scalability to larger image sets.
3D and video-based approaches. Distinct from these methods, 3D editing techniques [18, 11, 46, 47] enable manipulation of static multi-view sets through rendered views but operate on 3D representations rather than directly editing 2D images. This imposes restrictive assumptions that limit their applicability to image sets with unknown camera parameters, articulated objects, and varying backgrounds. Similarly, video editing and generation methods (e.g., [17, 52, 45]) are unsuitable for our task, as they assume temporal continuity and rely on motion coherence that does not hold across unordered image sets.
@@ 3 Method
We build on a pre-trained, frozen, depth-conditioned T2I model. Although designed for single-image generation, these models have been shown to produce image grids when prompted with joint layouts (e.g., “Side-by-side views of…”), establishing cross-image relationships as demonstrated in recent work [22, 38]. However, this emerged capability, which we refer to as the grid prior, exhibits several key limitations: (i) it provides only partial consistency in appearance, shape, and semantics (Fig. 2b); (ii) the consistency deteriorates rapidly as more images are composed (Fig. 2b); and (iii) generating a single canvas is bounded by the model’s native resolution, limiting scalability.
Our method leverages the grid prior while overcoming its core limitations (Fig. 2d). Specifically, we model the image set as a Pairwise Consistency Graph, comprising all possible two-image grid generations (Sec. 3.2). This allows us to exploit the strong inductive bias of the grid prior, while eliminating its scale limitation. To enhance visual consistency both within each image grid and across grids, we perform joint feature manipulation across all pairwise generations. To this end, we utilize dense 2D correspondences from the source set to automatically identify shared regions – without requiring object masks – and enforce fine-grained alignment. We find that feature-space similarity along these matches correlates strongly with visual coherence, motivating the use of Multiview Feature Fusion (Sec. 3.3). We further refine details via Feature Guidance, using a feature-matching objective (Sec. 3.4). Our full pipeline is illustrated in Fig. 3.
@@ 3.2 Pairwise Consistency Graph
Since every node $z^{t}_{i}$ participates in multiple adjacent edges, a graph denoising step yields multiple versions of $z^{t-1}_{i|ij}$. We thus consolidate pairwise latents $\{z^{t-1}_{ij}\}_{e\in E}$ back into image latents $\{z^{t-1}_{i}\}_{i=1}^{N}$ by extracting and averaging all $z^{t-1}_{i|ij}$, following [4]. In practice, full graph connectivity is not required, as analyzed in Sec. 4.4.
@@ 3.3 Multiview Feature Fusion
Matches between each image pair are defined as a partial feature coordinate mapping $M_{ij}:\mathcal{C}_{i}\!\to\!\mathcal{C}_{j}$, where a coordinate $\mathbf{c}\in\mathcal{C}_{i}$ corresponds to $\mathbf{c}^{\prime}=M_{ij}(\mathbf{c})\in\mathcal{C}_{j}$ and $\mathcal{C}_{i}$ contains all matched points (Fig. 4b). We first address pairwise consistency by considering a two-node graph with a single edge. Let $\mathbf{f}_{ij}\in\mathbb{R}^{H\times 2W\times D}$ be a feature map from a model’s forward pass on a grid, interpreted as $\mathbf{f}_{ij}=\operatorname{concat}(\mathbf{f}_{i},\mathbf{f}_{j})$, with $\mathbf{f}_{i},\mathbf{f}_{j}\in\mathbb{R}^{H\times W\times D}$. Given a pooling operator $\mathbf{f}[\mathbf{c}]$ extracting the vector at $\mathbf{c}$, Multiview Feature Fusion (MFF) fuses matched coordinates:
@@ 3.4 Feature Guidance
Figure 5: Matched feature similarity vs. visual consistency. We consider increasingly consistent generation (left to right): random images $\to$ adding descriptive prompts $\to$ adding control signals $\to$ generating in a grid $\to$ DDIM [40] inversion which reconstructs fully consistent source images. Keys and values differ in scale but follow the same pattern: cosine similarity at matched locations rises with consistency. Dashed lines show the baseline all-to-all similarity of feature maps. Points are averaged over 40 image pairs, all correspondences, blocks, and timesteps.
@@
To align small details, we further perform feature guidance [13, 9, 36], completing each generation step (Fig. 3, right). Specifically, we define an optimization objective over all edges as the distance between matched features, and refine all node latents via gradient descent with $\nabla_{z^{t-1}_{i}}L_{\mathrm{guide}}$ of:
@@ 4 Experiments
Figure 9: Qualitative comparisons. Match-and-Fuse (Ours) generates image sets with the highest consistency compared to our baselines FLUX [6], FLUX Kontext [25], IC-LoRA [22], and Edicho [3]. See Sec. 4.3 for comments on each result and SM for full image sets.
Figure 10: Extended Applications. Our method generalizes across various settings, enabling (a) consistent story generation from sketches and (b) localized editing omitting $\mathcal{P}^{theme}$ via FlowEdit [23] integration. See full sets of results in SM.
We extensively evaluate Match-and-Fuse. Our implementation builds on FLUX [6] with depth-conditioning [49], and uses RoMA [12] for matching. Full details are in SM.
@@ 4.1 Evaluation Setup
Evaluating our task requires image sets that share content across diverse scenarios, along with a metric that captures fine-grained cross-image consistency. Since no existing benchmark or metric supports this setting, we introduce a setup tailored to consistent set-to-set generation.
Evaluation set.  We curate a benchmark with 400 edits for 149 diverse distinct image sets, 3–15 images in each, combining all sets from [24] and [34] with frames sampled from 3D datasets [18, 5, 26], keyframes of a few publicly sourced videos, and sketched storyboards generated by ChatGPT. Data will be publicly released.
@@
Metrics.  We quantitatively evaluate our method, beginning with fine-grained cross-image consistency. Previous work relies on global visual-similarity metrics to assess set consistency—for example, [42] uses DreamSim [15] on object-masked regions. However, as our user study (Tab. 2b) shows, such global metrics are not ideal. To address this, we introduce DINO-MatchSim. For each image pair, we compute patch-level nearest-neighbor correspondences $\textit{NN}_{ij}$ between DINOv3 [39] feature maps from the source images, and measure similarity at the corresponding output locations:
@@
where $\tilde{F}_{i}$ denotes the $i^{\text{th}}$ output feature map. The average similarity across all image pairs is DINO-MatchSim. For prompt adherence, we report the average CLIP score over the set. To assess human preference, we conduct a large-scale user study on a random subset of the benchmark, comparing Match-and-Fuse to a baseline in a randomized two-alternative forced-choice (2AFC) format. Finally, we include a VLM-based evaluation using GPT-5 [31], which received the same content as the human raters.
@@ 4.3 Comparisons
Since no existing method is designed for our set-to-set generation task, we compare Match-and-Fuse to the closest baselines: (1) FLUX [6] with ControlNet [49] conditioning serves as a baseline for assessing independent generation using the same base model as our method. (2) FLUX Kontext [25], an image-to-image editing model. While it provides consistent edits over multiple turns on a single image, it lacks an explicit mechanism for achieving consistency across a single-turn image set editing. (3) IC-LoRA [22], text-driven generation of image sets with shared elements by fine-tuning LoRA modules for specific context types. We use their most relevant public checkpoint and add spatial ControlNet conditioning for comparability. (4) Edicho [3], the closest baseline to our method, combines depth-conditioned generation with explicit correspondences, but is limited to an image pair editing through a single reference. Additional implementation details are in Sec. A.2.
Figure 9 shows representative qualitative results. FLUX [6] produces inconsistencies in both coarse and fine details. FLUX Kontext [25] has low prompt adherence (e.g., a toy remains animal-like), and often distorts object structure (e.g. vase, dog). IC-LoRA [22] achieves partial consistency, with coherence often restricted to subsets (vase). Its realism and fidelity are notably lower than FLUX. Edicho [3] performs best among baselines but still shows noticeable inconsistencies (dog), as its one-to-all warping enforces consistency only with the first image, making the choice of an anchor ambiguous and degrading coherence across views. In contrast, Match-and-Fuse (bottom row) maintains high image quality while substantially improving structural and fine-grained consistency. We additionally compare to the closed-source Nano Banana API [10] and find that it does not solve our task (SM).
@@ 4.5 Extended Applications
As shown in Fig. 10a, our method generalizes to sketched source sets, enabling consistent visualization of storyboards. In Fig. 10b, we further demonstrate localized editing by combining our method with FlowEdit [23]. We provide default integration parameters in the SM; however, as is a known limitation of FlowEdit, achieving the desired balance between structure preservation and appearance change often requires mild per-edit hyperparameter tuning. Automating this selection is left for future work.
@@ A.1 Method
Preprocessing. Prompt construction with a VLM (Sec. 3.1) is used as an automated, scalable alternative to manual prompting. We instruct GPT-4o [30] via the system prompt (Fig. A.3) to generate $P^{non-shared}_{i}$ and enrich $\mathcal{P}^{\text{shared}}$ with creative artistic details (e.g., colors, textures). Figure A.1 exemplifies VLM input and output. Source 2D correspondences and shared content are identified using RoMA [12] predictions filtered by a confidence threshold $\operatorname{c}>0.05$.
General. All results are generated using FLUX [6] with $t=25\dots 0$ steps. We apply ControlNet [49] complementing DepthAnythingV2 [50] depth with HED edge maps [48] with equal weights, and linearly anneal control strength $w_{\text{ctrl}}^{t}=\text{linear\_anneal}(1,0,t)$. This enforces correct layout while preventing over-constraining control.
@@
Localized editing. FlowEdit [23] is an inversion-free method that enables background-preserving editing by traversing a path between source and target image distributions with $v$-predictions $V=V^{tar}(z^{tar})-V^{src}(z^{src})$. To incorporate it into our method, we redefine $V_{ij}$ implied in $\operatorname{denoise}_{\operatorname{MFF}}(\cdot)$ in Eq. 1 as $V_{ij}=V^{tar}_{ij}-V^{src}_{ij}$, where
@@
Similar to [28], FlowEdit relies on a single starting timestep parameter $n_{\text{max}}$ to balance structure preservation and appearance deviation. We ease this trade-off by adopting spatial control with an earlier start step. The used hyperparameters are: $T=25,n_{\text{max}}=24,n_{\text{min}}=10,w_{\text{ctrl}}=0.5,t^{MFF}_{\text{d}}>20$. Feature Guidance is omitted.